Displaying report 1-1 of 1.
Reports until 18:31, Wednesday 24 July 2024
Atoms General
Briana Chen - posted 18:31, Wednesday 24 July 2024 (11774)
EOM alignment, soldering

[Briana, Ian]

The overnight measurement of polarization drift was not a success, the channel setting needs to be 4 Vpp on the data acquisition since the voltage readings are not at the height of 40 Vpp. Redoing tonight. As for the polarization maintaining fiber, Thorlab people said you can't change the input of our butterfly laser (which would have been useful for ensuring the polarization matches the fiber's slow axis) and that the output should be along the slow axis. Maybe you actually can, but I'm worried about breaking things: the blue coating seems pretty attached to the butterfly pins. I'm not sure if rotating the laser output fiber will change anything, but should look at the drift to see if it is significant before attempting this. 

Aligned EOM. For an input of 6.6 mW, the output is around 6.5 mW f rom the photodetector, so it is reasonably aligned- the 5 axis stage is beautiful.

Soldered piezo today, which consisted of soldering to corresponding black/red wires to lengthen the smaller wires on the piezo. Then, the black wire was soldered to the silver pin on the back of a BNC connector and the red wire soldered to the gold pin. Twirled the wires around each other and learned to solder better by heating up the wire before putting the solder on. The tweezer solders are great for clearing out solder stuck in holes. Good preparation for soldering the SMA to the BNC cable for the EOM. 

Tomorrow, I'll redo the dip depth measurements. 

Images attached to this report
Displaying report 1-1 of 1.